In cases where an employee is paid different rates for different jobs during the same workweek, the employers can, at their option, calculate the overtime in one of three ways: Higher Rate for Overtime Hours, Based on Actual Hours Worked, or Rate after the 40th Hour.

Using the hours and rates for Eli Employee, Example's son, his gross wages are computed below using the three different methods.

Eli makes $14.75 per hour for Job A and $19.05 per hour for Job B and of the 44 hours worked during the week, he completed 20 hours for Job A, switched and completed 14 hours for Job B and finished up the week with 10 hours back on Job A.

(1) Higher Rate for Overtime Hours

Under this method, the higher (or highest) rate paid is used to determine the overtime rate (highest rate times .5). This rate is then multiplied by the overtime hours, resulting in the overtime wages.

  1. Hours Formula Amount
    Hours Worked, Job A: 30.0 30.0 hours × $14.75 = $442.50
    Hours Worked, Job B: 14.0 14.0 hours × $19.05 = 266.70
    Overtime hourly rate ($19.05a × 0.5) = $9.53, rounded
    Overtime Hours: 4.0 $9.53 × 4.0 hours = 38.12
    Gross Earnings = $747.32
    a = the higher of the two

(2) Based on Actual Hours Worked

Under this method, the overtime rate is computed (total earnings divided by total hours worked multiplied by .5). This rate is then multiplied by the overtime hours, resulting in the overtime wages.

  1. Hours Formula Amount
    Hours Worked, Job A: 30.0 30.0 hours × $14.75 = $442.50
    Hours Worked, Job B: 14.0 14.0 hours × $19.05 = 266.70
    Overtime hourly rate $709.20a ÷ 44.0 hours = $16.12, rounded
    $16.12 × 0.5 = $8.06, rounded
    Overtime Hours: 4.0 $8.06 × 4.0 hours = 32.24
    Gross Earnings = $741.44
    a = combine wages for both Job A and Job B

(3) Rate after the 40th Hour

Under this method, the overtime time rate for the job performed during overtime hours is used (overtime hour rate multiplied by .5). This rate is then multiplied by the overtime hours, resulting in the overtime wages.

  1. Hours Formula Amount
    Hours Worked, Job A: 30.0 30.0 hours × $14.75 = $442.50
    Hours Worked, Job B: 14.0 14.0 hours × $19.05 = 266.70
    Overtime hourly rate ($14.75a × .5, rounded) × 4.0 hours = 29.52
    Gross Earnings = $738.72
    a = the rate ($14.75, Job A) after the 40th hour

? Tackle It

Note: Round interim calculations to two decimal places. Use rounded amounts in subsequent computations. And, round your final answers to two decimal places.

Compute Alex Rodriguez's earnings using the three methods described above. Assume the rate for Job A is $17.05 per hour and $18.45 per hour for Job B. Of the 43 hours worked during the week, 25 hours were completed for Job A and 18 hours complete for Job B.

Higher Rate for Overtime Hours

Regular Hours, Job A Hourly Rate, Job A $ Regular Total Pay, Job A $
Regular Hours, Job B Hourly Rate, Job B Regular Total Pay, Job B   
Overtime Hours Hourly Rate, Overtime Overtime Total Pay   
Total Gross Earnings $

Based on Actual Hours Worked

Regular Hours, Job A Hourly Rate, Job A $ Regular Total Pay, Job A $
Regular Hours, Job B Hourly Rate, Job B Regular Total Pay, Job B   
Overtime Hours Hourly Rate, Overtime Overtime Total Pay   
Total Gross Earnings $

Rate after 40th Hour

In this scenario, Alex worked a total of 48 hours, the regular hours evenly spread between Job A (at a rate of $15.85/hour) and Job B (at a rate of $18.45/hour). The overtime hours can be attributed to Job B.

Regular Hours, Job A Hourly Rate, Job A $ Regular Total Pay, Job A $
Regular Hours, Job B Hourly Rate, Job B Regular Total Pay, Job B   
Overtime Hours Hourly Rate, Overtime Overtime Total Pay   
Total Gross Earnings $

? Medal Points

Of the three methods described above, one requires that the employee agree to the method in advance. It is the   method.
Answer #1

Solution:

The information given is as follows:

The rate for Job A = $17.05 per hour

The rate for Job B = $18.45 per hour

Hours completed for Job A = 25 hours

Hours completed for Job B = 18 hours

Total hours worked = 43 hours

Overtime hours= 43-40 = 3 hours

  1. Higher Rate for Overtime Hours

Under this method, the higher (or highest) rate paid is used to determine the overtime rate (highest rate times .5). This rate is then multiplied by the overtime hours, resulting in the overtime wages.

Hours

Formula

Amount

Hours Worked, Job A: 25.0

25.0 hours × $17.05

=

$426.25

Hours Worked, Job B: 18.0

18.0 hours × $18.45

=

$332.10

Overtime hourly rate

($18.45a × 0.5) = $9.23, rounded

Overtime Hours: 3.0

$9.23 × 3.0 hours

=

$22.69

Gross Earnings

=

$786.04

a = the higher of the two

  1. Based on Actual Hours Worked

Under this method, the overtime rate is computed (total earnings divided by total hours worked multiplied by .5). This rate is then multiplied by the overtime hours, resulting in the overtime wages.

Hours

Formula

Amount

Hours Worked, Job A: 25.0

25.0 hours × $17.05

=

$426.25

Hours Worked, Job B: 18.0

18.0 hours × $18.45

=

$332.10

Overtime hourly rate

($758.35a × 43.0 hours) = $17.64, rounded

($17.64 × 0.5) = $8.82, rounded

Overtime Hours: 3.0

$8.82 × 3.0 hours

=

$26.45

Gross Earnings

=

$784.80

a = combine wages for both Job A and Job B

  1. Rate after the 40th Hour

Under this method, the overtime time rate for the job performed during overtime hours is used (overtime hour rate multiplied by .5). This rate is then multiplied by the overtime hours, resulting in the overtime wages.

The information given is as follows:

The rate for Job A = $15.85 per hour

The rate for Job B = $18.45 per hour

Total hours worked = 48 hours

Hours completed for Job A = 20 hours

Hours completed for Job B = 48-20 = 28 hours

Overtime hours= 48-40 = 8 hours

Solution:

Hours

Formula

Amount

Hours Worked, Job A: 25.0

20.0 hours × $15.85

=

$317.00

Hours Worked, Job B: 18.0

28.0 hours × $18.45

=

$516.60

Overtime hourly rate

($18.45a × 0.5 rounded) * 8.0 hours

=

73.8

Gross Earnings

=

$907.40

a = the rate ($18.45, Job B) after the 40th hour

Hence, The Gross earnings is the highest in method 3 with rate after 40th hour which is $907.40.